OpenStudy (taylev105):

f What is the simplified expression for 5 to the power of negative 3 multiplied by 2 to the power of 2 multiplied by 5 to the power of 6 whole over 2 to the power of 3 multiplied by 5 to the power of 2?

OpenStudy (taylev105):

lol please help again @dumbcow

OpenStudy (dumbcow):

lol post a pic ... i cant read that

OpenStudy (dumbcow):

why dont i let you solve this one here are the properties you need to know: \[x^m * x^n = x^{m+n}\] \[\frac{x^m}{x^n} = x^{m-n}\]

OpenStudy (taylev105):

???????

OpenStudy (dumbcow):

do you not understand the properties of exponents? all you are doing is adding or subtracting the exponents for each base (2 or 5)

OpenStudy (taylev105):

lolol im sooo confused

OpenStudy (dumbcow):

haha ok try to do it one step at a time \[5^{-3} * 5^6 = ?\]

OpenStudy (taylev105):

125

OpenStudy (taylev105):

@dumbcow

OpenStudy (dumbcow):

sorry yes thats right leave it as 5^3 for now though next step.... what is \[\frac{5^3}{5^2}\]

OpenStudy (dumbcow):

evaluate \[\frac{2^2}{2^3}\] in the same way then you will have your answer :)`

OpenStudy (taylev105):

I got 125/25 and 4/8 whats the sanswer ???

OpenStudy (taylev105):

@midhun.madhu1987

OpenStudy (taylev105):

is 5/2 correct ???? plesase help!!! will fan and medal @Mvpriest6 @ugh.math

OpenStudy (taylev105):

@AriPotta

OpenStudy (taylev105):

@HelpBlahBlahBlah

OpenStudy (taylev105):

@srgntb0$$

OpenStudy (taylev105):

is 5/2 correct?

OpenStudy (taylev105):

@karatechopper

OpenStudy (taylev105):

can someone help me please....................................

OpenStudy (taylev105):

please tell me its right??

OpenStudy (taylev105):

@geerky42 help me please??

OpenStudy (taylev105):

@silverr you are my only chance

OpenStudy (taylev105):

is 5/ 2 correct?

OpenStudy (anonymous):

let me look, okay? i might be able to help i'm your only chance omg

OpenStudy (anonymous):

computer glitch hurr hurr. it's simple PEMDAS. 5^-3 = .008 2^2 = 4 5^6 = 15625 multiply all those and you get 500. one second and i find the rest.

OpenStudy (anonymous):

2^3 = 8 5^2 = 25 25 * 8 = 200 500 over 200 both divided by 100 = 5 over 2, or 5/2 so, you are correct ! and if you want a whole #, it's 2 1/2.

OpenStudy (anonymous):

@Taylev105 ? <33

OpenStudy (taylev105):

no it has to be a equation...

OpenStudy (anonymous):

what do you mean ?

OpenStudy (taylev105):

the answer has to be a equation..

OpenStudy (anonymous):

are there answer choices? i don't get what you mean.

OpenStudy (taylev105):

2 to the power of 2 over 5 to the power of 2, 5 over 2 to the power of 2, 5 over 2, 5 to the power of 2 over 2

OpenStudy (anonymous):

OH I SEE. one second. i will see if i can convert or something.

OpenStudy (taylev105):

ok

OpenStudy (taylev105):

@KamiBug can you help us ???

OpenStudy (anonymous):

it's 5/2.

OpenStudy (taylev105):

ok thank you

OpenStudy (anonymous):

no problem :)